Solve the following using the elimination method.
5y - 3x = -20
3x - 3y = 14
I'm having trouble with these methods. So if you could show me how you got the numbers you got and what you did exactly, that'd be great.

Answer by atique.ah(23)

Solve the following using the elimination method.
5y - 3x = -20..........equation (1)
3x - 3y = 14...........equation (2)
solution:
first arrange te equation in the standard form(ax+by=c), ie x then y = and constant

-3x + 5y = -20 (arranging equation 1)
3x - 3y = 14 (equation 2)

if you look at the two equations carefully you will notice there is -3x in equation 1 and 3x in equation 2 so if we add the two equations 3x and -3x will cancell each oter and we will be left with only "y". so let us do it
5y - 3y = -20 +14
2y = -6
y = -3
substitute this value in any equation and get te value of y. let us use equation 2
3x - 3(-3)=14
3x + 9 = 14
3x = 14 - 9
3x = 5
x = 5/3
Answer x = 5/3 and y = -3